OpenStudy (anonymous):

7^4/7^2 ____ 49 Replace the ___ with <, >, or = to make a true statement.

OpenStudy (anonymous):

\[\frac{ 7^4 }{ 7^2 }=7^{4-2} \ [ \ ? \ ] \ 49\] Which symbol should replace the question mark? @yolifuentes67

OpenStudy (texaschic101):

7^4 = 2401 7^2 = 49 2401/49 = 49 49 ( ? ) 49 what symbol goes where the question mark is ?